OpenStudy (anonymous):

Just my weekly earn one medal question! (first to answer wins) (Getting a little harder, but it is still easy!) xs+d=t What is "x?"

OpenStudy (imstuck):

What do you mean by "what is x"?

OpenStudy (anonymous):

I mean find the value of x.

OpenStudy (texaschic101):

xs + d = t --- subtract d from both sides xs = t - d --- divide by s x = (t - d)/s or x = t/s - d/s

OpenStudy (anonymous):

Correct, nice job. Question closed.

OpenStudy (texaschic101):

lol....thanks